OpenStudy (anonymous):

Someone please help me. Really appreciated. Find x if y = -4 3y - 6/x = 12 A = 0 B = -1/4 C = 6 D = -4

OpenStudy (harsimran_hs4):

y = -4 so subsitute this value in the given equation and solve for x ...can you ?

OpenStudy (anonymous):

Yes, but I get lost. @harsimran_hs4

OpenStudy (harsimran_hs4):

ok 3y - 6/x = 12 putting value of y 3(-4) - 6/x = 12 -12 -6/x = 12 6/x = -24 can you proceed now ?

OpenStudy (anonymous):

-1/4 :)

OpenStudy (harsimran_hs4):

yes

OpenStudy (anonymous):

Thankkkkkkyouuuuuu. <3 :)